# Superconducting spintronics with electron symmetry filtering and interfacial spin-orbit coupling

### v1

@@ III Junction properties in the normal state
We begin with a discussion of the junctions normal-state properties, including their non-volatile magnetization configurations and tunneling magnetoresistance, with special emphasis on the surprising TMR enhancement under applied bias observed in S2 junctions.
It is well known Hallal2013 that interfacial SOC provides a perpendicular magnetic anisotropy (PMA) in MgO/Fe structures, responsible for the perpendicular magnetization alignment in thin (below 2 nm) films Ikeda2010. For thicker Fe films, PMA is expected to compete with in-plane magnetic anisotropy. As a consequence, for the most part of this work we selected a 10 nm thick magnetically free Fe layer, which enables three orthogonal and non-volatile magnetic states (two in-plane and one out-of-plane) without requiring an external magnetic field. As we show in the next chapters, these configurations allow us to systematically study the influence of the magnetization direction on the response of S/F hybrids (without the application of any external magnetic field), experiments that were previously unfeasible in superconducting spintronics.
@@ III.1 Non-volatile magnetization states
In our junctions, the different non-volatile magnetic configurations are achieved by applying and removing an external magnetic field using a 3D vector magnet composed of three superconducting coils. The three axes of the magnet are aligned with the main crystallographic directions of the junctions, and the applied field can reach 1 T in two of the axes and 3.5 T in the other one. In the N/F/F MTJs (S2 in Figure 1), one of the electrodes is made to be a soft F (10 nm Fe) and the other a hard F (Fe/Co). The relative alignment of the two F layers is found by measuring the conductance at low biases (<10 mV). The magnetization’s behavior (remanence, coercive fields, etc) depends on the competition between crystalline, volume (or shape) and surface anisotropies, as well as the interaction between both F layers. Due to the growth of the samples, the plane of the layers and the crystalline Fe bcc easy axis are the same (100)Tiusan-2007. Because of that, in addition to the shape anisotropy, the coercive fields of the F layers are lower in the in-plane directions than in the out-of-plane one, and thereby the F magnetization will follow much more easily the external magnetic field if it is applied in-plane.
@@
Figure 3: (a,b) In-plane TMR in the easy axis [100] for a a N/F/F junction. Both parallel and anti-parallel magnetic alignments are observed, with a very low coercive field ($\sim 25-50$ Oe) for the soft layer and a higher ($\sim 400-600$ Oe) for the hard one. (c,d) In-plane rotation of the magnetic field, using a field higher than the soft layer coercive field (80 Oe), making it reorient following the field and MCA, but lower than the hard layer coercive field, so it is fixed during the experiment. (e,f) Out-of-plane TMR at $T=5$ K with an OOP alignment of the soft layer. Panels (a,c,e) sketch the orientation of the magnetic field with respect to the MCA in each of the experimetns. Figure adapted from refs. GonzalezRuano2020; GonzalezRuano2021 with the necessary permissions.
@@
We already mentioned that in order to study the interaction between S and F, it is important to have different remanent magnetic configurations available. Figure 3a shows a TMR experiment in the N/F/F sample (S2) where $H$ is swept along the Fe [100] crystallographic direction. It reveals the existence of a parallel (P) and antiparallel (AP) in-plane states of the two ferromagnetic electrodes. The coercive fields of the soft and hard magnetic layers can also be extracted, being 25-50 Oe and 400-600 Oe, respectively. The in-plane rotation sketched in Figure 3b shows the three possible in-plane magnetic configurations: The above mentioned P and AP states and the perpendicular in-plane (PIP) state. By making a rotation with an external field lower than the hard F coercive field, we maintained that magnetization fixed. The soft F magnetization follows the external field in jumps between four different directions which correspond to the direction of the Fe crystallographic easy axis. The lower effective TMR (around 40 $\%$) value is not due to the crystalline structure’s quality, but arises from the in-series configuration of the N/F/F junction which includes a normal metal electrode (vanadium), providing a nearly constant resistance which limits the overall TMR ratio.
@@
Finally, Figure 3c shows an experiment similar to that in panel (a) but with $H$ applied perpendicularly to the plane of the sample. It can be deduced that there is an additional out-of-plane (OOP) magnetization state in which the hard layer remains magnetized IP and the soft layer is OOP. The transition from the P to the OOP state occurs at $H_{OOP}\simeq 1.5$ kOe. Importantly, this OOP state is remanent, i.e. it is persistent when $H_{OOP}$ is removed, and a magnetic field of 400 Oe in the opposite direction is needed to switch back to the P in-plane state. Recent trends in spintronics using MTJs take advantage of PMA to provide large perpendicular room temperature TMR, enhanced thermal stabilityIkeda2010, low spin-torque switching currentsLeutenantsmeyer2015; Lau2016 and record small lateral sizesIgarashi2017. As we have shown above, our 10 nm thick magnetically free Fe layers (due to competing crystalline, shape and surface anisotropies) enable us to change the magnetization of the film either IP or OOP so that the magnetization stays remanent at low temperatures when switching off the external field, as it is shown in figure 3c.
@@
To elucidate the mechanism behind the different magnetization orientations due to competing anisotropies we carried out micromagnetic simulations using MuMax3mumax, considering three models that differ in their in-depth anisotropy. The simulations described in Ref.  Martinez2018 show that the spin-flip transition is rather weakly affected by the presence of the Fe layer cubic anisotropy. This demonstrates that it is mainly a demagnetization energy competing with PMA which provides the spin reorientation transition.
@@ III.2 Ab-initio SOC calculations at the V/MgO and Fe/MgO interfaces
To further understand our system, the theoretical study of the Rashba SOC and its electric field dependence has been performed using the ab-initio Full Potential Linear Augmented Plane Wave FP-LAPW code WIEN2k Wien2k within a fully relativistic spin-orbit scheme. In our calculations, we used a supercell model thoroughly chosen to describe the X/MgO(001) interface where X = V and Fe (see Figure 4a). The Brillouin zone was sampled using a $25\times 25\times 1$ $k$-point mesh, which was selected based on a preliminary convergence study of the total energy with respect to the number of $k$-points. Given the strong sensitivity of such calculations to the energy convergence parameters in WIEN2k, particularly when relativistic local orbitals (LAPW+lo) are included, we employed $E_{max}=100$ Ry where $E_{max}$ is the maximum energy cutoff for plane waves, controlling the size and accuracy of the basis set. This choice ensures that all scalar-relativistic eigenstates are taken into account when SOC is enabled. The basis set size for the wavefunction expansion was chosen as $R{\times}K_{max}=7$, where $R$ is the smallest muffin-tin radius, $RMT$, and $K_{max}$ is the largest reciprocal lattice vector. Following the zig-zag potential electric field implementation in the WIEN2k code introduced by Stahn et al. Stahn2001, the supercell is constructed such that one half contains the multilayer sequence, while the other half is a vacuum region. This setup generates a constant electric field, $E=-\Delta{V}/\Delta{z}$, localized at the X/MgO interface. Both the amplitude and the sign of the field are determined by the applied potential ramp and whether the ramp is ascending or descending within that region. Therefore, for different values of the electric field, we calculated the Rashba parameter $\alpha_{R}$ from the band structure obtained from WIEN2k.
@@
Figure 4: (a) Supercell model for the X/MgO interfaces with X = V, Fe. Calculated band splitting due to Rashba SOC at the (b) V/MgO and (c) Fe/MgO interface in the absence of electric field. Inset of (b): variation of $\alpha_{R}$ with the electric field in case of V/MgO interface. Charge density at the X/MgO interface (blue rectangle in (a)) for (d) V/MgO and (e) Fe/MgO. (f) Sketch model for the Rashba splitting of the parabolic bands, the $k$ lines in (b) and (c) correspond to $k_{y}=0$, the direction in the k space (100) corresponds to $k_{x}>0$ and (-100) to $k_{x}<0$.
@@
The spin-orbit term of the non-relativistic Dirac (Pauli) Hamiltonian: $H_{SO}=\nicefrac{{\hbar}}{{(2m_{0}c)^{2}}}\overrightarrow{\nabla}V\cdot(\overrightarrow{\sigma}\times\overrightarrow{p})$ can be implified for 2 dimensional electronic systems with the confinement direction (e.g. $\overrightarrow{OZ}$ in a Cartesian system) perpendicular to the propagation direction (within the $XOY$ plane) leading to the Rashba Hamiltonian: $H_{R}=\alpha_{R}\overrightarrow{\sigma}\cdot(\overrightarrow{k}\times\overrightarrow{e_{z}})$ where $\alpha_{R}=\nicefrac{{\hbar^{2}}}{{(2m_{o}c)^{2}}}{\partial{V}}/{\partial{z}}$ is the Rashba constant which is a measure of the spin-orbit interaction and $\overrightarrow{e_{z}}$ is the unit vector of the $\overrightarrow{OZ}$ (electron confinement) direction. From its definition, $\alpha_{R}$ is proportional to the electric field $E=-{\partial{V}}/{\partial{z}}$ aligned along the $\overrightarrow{OZ}$ direction. At the interface between two dissimilar materials, interfacial electric fields naturally arise, giving rise to Rashba-type spin–orbit interaction effects. The strength of the SOC is further enhanced at metal surfaces, where the breaking of translational symmetry is equivalent to a potential gradient experienced by the electrons. In multilayer heterostructures comprising a metal–insulator or metal–semiconductor interface, the formation of a depletion region further induces a pronounced interfacial electric field. By diagonalizing the Rashba Hamiltonian, the eigenvalues will be: $E_{\pm}(k_{\parallel})=\nicefrac{{\hbar^{2}k_{\parallel}^{2}}}{{2m_{0}}}\pm\alpha_{R}|{k}_{\parallel}|$ representing parabolic bands with an offset of the parabola minimum in positive (or negative) $k$ values (Rashba splitting). The minimum of the parabola can be found by ${\partial{E}}/{\partial{k}}=0\Rightarrow{k}_{0}={m_{0}\alpha_{R}}/{\hbar^{2}}$ so that $E_{min}=E_{R}=\nicefrac{{\hbar^{2}k_{0}^{2}}}{{2m_{0}}}$ that leads to $E_{R}={k_{0}\alpha_{R}}/{2}$. From this equation, $\alpha_{R}$ can be calculated from the values of $E$ and $k$ corresponding to the minimum of the parabolic dispersion band: $\alpha_{R}={2E_{R}}/{k_{0}}$. We used this equation to extract the Rashba parameter from the Rashba offset of the parabolic bands calculated for the X/MgO/vacuum supercell models. Note that, experimentally, the SOC constant $\alpha_{R}$ can be extracted from Angular Resolved Photoemission (ARPES) Nicolay2001; Reinert2001.
@@ III.3 SOC-induced conductance bottleneck
The selection of epitaxial V(100)/MgO/Fe(100) junctions to investigate the SOC-induced interaction between a ferromagnet and a superconductor may initially seem counterintuitive. This is because the orbital symmetries in the electronic structures of V(100) and Fe(100) are mismatched, implying that heterojunctions based on these materials should exhibit negligible conduction under low-bias conditions. Indeed, early studies on fully epitaxial lateral bcc Fe/V/Fe spin valves have already highlighted the critical role of orbital symmetries in spin-dependent transport across a V/Fe interface Moodera2008.
At the Fermi energy $E_{F}$, V(100) possesses $\Delta_{2}$ orbital symmetry, whereas Fe(100) is dominated by $\Delta_{1}$ symmetry (see Figure 5a). Therefore, MgO acts as an insulating layer that suppresses $\Delta_{2}$ states and enables giant tunneling magnetoresistance (TMR) in Fe(100)/MgO(100)/Fe(100) structures Parkin2004; Yuasa2004; Butler2001; Mathon2001. Nevertheless, the structural inversion asymmetry in our junctions introduces interfacial SOC at the V/MgO interface, which works in conjunction with the effective $\Delta_{2}$ filtering barrier of crystalline MgO. As sketched in Figure 5a and discussed in Ref. GonzalezRuano2021-HB-TMR , SOC-driven spin-flip scattering mixes $\Delta_{2}$ and $\Delta_{1}$ states, opening a pathway for electron tunneling above $T_{C}$ at low bias, as depicted in Figure 5a.
Another channel contributing (in sequence) to the low bias transport in these epitaxial junctions originates from so-called “hot spots” in the reciprocal space provided by $\Delta_{1}$ states of electrons that undergo normal incidence to the barrier, allowing a strong transmission through MgO Butler2001; Mathon2001. The normal-state transport in this system can therefore be approximated by an effective resistance $R_{\text{eq}}=R_{\text{SOC}}+R_{\text{MgO}}$, where the resistance associated with the symmetry-enforced SOC barrier, $R_{\text{SOC}}$, is significantly higher than that the one arising from the conventional MgO barrier, $R_{\text{MgO}}$.
This framework, in which tunneling is governed by symmetry-based spin filtering rather than the barrier height, is supported by our measurements of the low-bias differential conductance G (Figure 5b). These measurements highlight the crucial role of SOC in electron/spin transport through the V/MgO boundary. Compared with junctions lacking such an interface, G is reduced by roughly two (three) orders of magnitude in the presence of one (two) interfaces. This highlights a unique situation in (superconducting) spintronics in which practically all electrons that are transmitted between V and Fe undergo the effects of SOC.
Figure 5: (a) Transport mechanism in the V/MgO/Fe barrier, sketching the vanadium mixed $\Delta_{2}$ and $\Delta_{1}$ surface states, interfacial SOC and symmetry filtering in the MgO. (b) average conductance for different types of samples of the same lateral size ($20\times 20~\mu\text{m}^{2}$), as a function of the number of V/MgO interfaces for each type.

### v2

@@ III Junction properties in the normal state
We begin with a discussion of the junctions normal-state properties, including their non-volatile magnetization configurations and tunneling magnetoresistance, with special emphasis on the surprising TMR enhancement under applied bias observed in the S2 type junctions.
It is well known Hallal2013 that the interfacial SOC provides a perpendicular magnetic anisotropy (PMA) in MgO/Fe structures, responsible for the perpendicular magnetization alignment in thin (below 2 nm) films Ikeda2010. For thicker Fe films, PMA is expected to compete with in-plane magnetic anisotropy. As a consequence, for the most part of this work, we selected a 10 nm thick magnetically free Fe layer, which enables three orthogonal and non-volatile magnetic states (two in-plane and one out-of-plane) without requiring an external magnetic field. As we show in the next chapters, these configurations allow us to systematically study the influence of the magnetization direction on the response of S/F hybrids (without the application of any external magnetic field), experiments that were previously unfeasible in superconducting spintronics.
@@ III.1 Non-volatile magnetization states
In our junctions, the different non-volatile magnetic configurations are achieved by applying and removing an external magnetic field, using a 3D vector magnet composed of three superconducting coils. The three axes of the magnet are aligned with the main crystallographic directions of the junctions, and the applied field can reach 1 T in two of the axes and 3.5 T in the other one. In the N/F/F MTJs (S2 in Figure 1), one of the electrodes is made to be a soft F (10 nm Fe) and the other a hard F (Fe/Co). The relative alignment of the two F layers is found by measuring the conductance at low biases (<10 mV). The magnetization’s behavior (remanence, coercive fields, etc) depends on the competition between crystalline, volume (or shape) and surface anisotropies, as well as the interaction between both F layers. Due to the growth of the samples, the plane of the layers and the crystalline Fe bcc easy axis are the same (100)Tiusan-2007. Because of that, in addition to the shape anisotropy, the coercive fields of the F layers are lower in the in-plane directions than in the out-of-plane one, and thereby the F magnetization will follow much more easily the external magnetic field if it is applied in-plane.
@@
Figure 3: (a,b) In-plane TMR in the easy axis [100] for a a N/F/F junction. Both parallel and anti-parallel magnetic alignments are observed, with a very low coercive field ($\sim 25-50$ Oe) for the soft layer and a higher ($\sim 400-600$ Oe) for the hard one. (c,d) In-plane rotation of the magnetic field, using a field higher than the soft layer coercive field (80 Oe), making it reorient following the field and magnetocrystalline anistropy (MCA), but lower than the hard layer coercive field, so it is fixed during the experiment. (e,f) Out-of-plane TMR at $T=5$ K with an out-of-plane alignment of the soft layer. Panels (a,c,e) sketch the orientation of the magnetic field with respect to the MCA in each of the experimetns. Figure adapted from refs. GonzalezRuano2020; GonzalezRuano2021 with the necessary permissions.
We already mentioned that in order to study the interaction between S and F, it is important to have different remanent magnetic configurations available. Figure 3a shows a TMR experiment in the N/F/F sample (S2) where the magnetic field ($H$) is swept along the Fe [100] crystallographic direction, where bias-dependent TMR is defined in terms of the relevant conductance, $G$ as function of an applied bias, $V$:
$\text{TMR}(V)=\dfrac{G_{P}(V)-G_{AP}(V)}{G_{AP}(V)}.$ (1)
@@
It reveals the existence of a parallel (P) and antiparallel (AP) in-plane states of the two ferromagnetic electrodes. The coercive fields of the soft and hard magnetic layers can also be extracted, being 25-50 Oe and 400-600 Oe, respectively. The in-plane rotation sketched in Figure 3b shows the three possible in-plane magnetic configurations: the above mentioned P and AP states and the perpendicular in-plane (PIP) state. By making a rotation with an external field lower than the hard F coercive field, we maintained that magnetization fixed. The soft F magnetization follows the external field in jumps between four different directions which correspond to the direction of the Fe crystallographic easy axis. The lower effective TMR (around 40 $\%$) value is not due to the crystalline structure’s quality, but arises from the in-series configuration of the N/F/F junction which includes a normal metal electrode (vanadium), providing a nearly constant resistance which limits the overall TMR ratio.
@@
Figure 3c shows an experiment similar to that in panel (a) but with $H$ applied perpendicularly to the plane of the sample. It can be deduced that there is an additional out-of-plane (OOP) magnetization state in which the hard layer remains magnetized IP and the soft layer is OOP. The transition from the P to the OOP state occurs at $H_{OOP}\simeq 1.5$ kOe. Importantly, this OOP state is remanent: it is persistent when $H_{OOP}$ is removed, and a magnetic field of 400 Oe in the opposite direction is needed to switch back to the P in-plane state. Recent trends in spintronics using MTJs take advantage of PMA to provide large perpendicular room temperature TMR, enhanced thermal stabilityIkeda2010, low spin-torque switching currentsLeutenantsmeyer2015; Lau2016, and record small lateral sizes Igarashi2017. As we have shown above, our 10 nm thick magnetically free Fe layers (due to competing crystalline, shape and surface anisotropies) enable us to change the magnetization of the film either IP or OOP so that the magnetization stays remanent at low temperatures when switching off the external field, as it is shown in Figure 3c.
@@
To elucidate the mechanism behind the different magnetization orientations due to competing anisotropies we carried out micromagnetic simulations MuMax3mumax, considering three models that differ in their in-depth anisotropy. The simulations described in Ref.Martinez2018 show that the spin-flip transition is rather weakly affected by the presence of the Fe layer cubic anisotropy. This demonstrates that it is mainly a demagnetization energy competing with PMA which provides the spin reorientation transition.
@@ III.2 Ab-initio SOC calculations at the V/MgO and Fe/MgO interfaces
To further understand the underlying electronic structure of our system, the theoretical study of the Rashba SOC and its electric field dependence has been performed using the ab-initio full potential linear sugmented plane wave FP-LAPW code WIEN2k Wien2k within a fully relativistic spin-orbit scheme. In our calculations, we used a supercell model thoroughly chosen to describe the X/MgO(001) interface, where X = V and Fe (see Figure 4a). The Brillouin zone was sampled using a $25\times 25\times 1$ $k$-point mesh, which was selected based on a preliminary convergence study of the total energy with respect to the number of $k$-points. Given the strong sensitivity of such calculations to the energy convergence parameters in WIEN2k, particularly when relativistic local orbitals (LAPW+lo) are included, we employed $E_{max}=100$ Ry where $E_{max}$ is the maximum energy cutoff for plane waves, controlling the size and accuracy of the basis set. This choice ensures that all scalar-relativistic eigenstates are taken into account when SOC is enabled. The basis set size for the wave function expansion was chosen as $R{\times}K_{max}=7$, where $R$ is the smallest muffin-tin radius, $RMT$, and $K_{max}$ is the largest reciprocal lattice vector. Following the zig-zag potential electric field implementation in the WIEN2k code introduced by Stahn et al. Stahn2001, the supercell is constructed such that one half contains the multilayer sequence, while the other half is a vacuum region. This setup generates a constant electric field, $E=-\Delta{V}/\Delta{z}$, localized at the X/MgO interface. Both the amplitude and the sign of the field are determined by the applied potential ramp and whether the ramp is ascending or descending within that region. Therefore, for different values of the electric field, we calculated the Rashba parameter $\alpha_{R}$ from the band structure obtained from WIEN2k.
@@
Figure 4: (a) Supercell model for the X/MgO interfaces with X = V, Fe. Calculated band splitting due to Rashba spin-orbit coupling at the (b) V/MgO and (c) Fe/MgO interface in the absence of electric field. Inset of (b): variation of $\alpha_{R}$ with the electric field in case of V/MgO interface. Charge density at the X/MgO interface (blue rectangle in (a)) for (d) V/MgO and (e) Fe/MgO. (f) Sketch model for the Rashba splitting of the parabolic bands, the $k$ lines in (b) and (c) correspond to $k_{y}=0$, the direction in the k space (100) corresponds to $k_{x}>0$ and (-100) to $k_{x}<0$.
@@
The spin-orbit term of the non-relativistic Dirac (Pauli) Hamiltonian: $H_{SO}=\nicefrac{{\hbar}}{{(2m_{0}c)^{2}}}\overrightarrow{\nabla}V\cdot(\overrightarrow{\sigma}\times\overrightarrow{p})$ can be implified for 2 dimensional electronic systems with the confinement direction (e.g., $\overrightarrow{OZ}$ in a Cartesian system) perpendicular to the propagation direction (within the $XOY$ plane) leading to the Rashba Hamiltonian: $H_{R}=\alpha_{R}\overrightarrow{\sigma}\cdot(\overrightarrow{k}\times\overrightarrow{e_{z}})$ where $\alpha_{R}=\nicefrac{{\hbar^{2}}}{{(2m_{o}c)^{2}}}{\partial{V}}/{\partial{z}}$ is the Rashba constant which is a measure of the SOC strength and $\overrightarrow{e_{z}}$ is the unit vector of the $\overrightarrow{OZ}$ (electron confinement) direction. From its definition, $\alpha_{R}$ is proportional to the electric field $E=-{\partial{V}}/{\partial{z}}$ aligned along the $\overrightarrow{OZ}$ direction. At the interface between two dissimilar materials, interfacial electric fields naturally arise, giving rise to Rashba-type SOC effects. The strength of the SOC is further enhanced at metal surfaces, where the breaking of translational symmetry is equivalent to a potential gradient experienced by the electrons. In multilayer heterostructures comprising a metal–insulator or metal–semiconductor interface, the formation of a depletion region further induces a pronounced interfacial electric field. By diagonalizing the Rashba Hamiltonian, the eigenvalues will be: $E_{\pm}(k_{\parallel})=\nicefrac{{\hbar^{2}k_{\parallel}^{2}}}{{2m_{0}}}\pm\alpha_{R}|{k}_{\parallel}|$ representing parabolic bands with an offset of the parabola minimum in positive (or negative) $k$ values (Rashba splitting). The minimum of the parabola can be found by ${\partial{E}}/{\partial{k}}=0\Rightarrow{k}_{0}={m_{0}\alpha_{R}}/{\hbar^{2}}$ so that $E_{min}=E_{R}=\nicefrac{{\hbar^{2}k_{0}^{2}}}{{2m_{0}}}$ that leads to $E_{R}={k_{0}\alpha_{R}}/{2}$. From this equation, $\alpha_{R}$ can be calculated from the values of $E$ and $k$ corresponding to the minimum of the parabolic dispersion band: $\alpha_{R}={2E_{R}}/{k_{0}}$. We used this equation to extract the Rashba parameter from the Rashba offset of the parabolic bands calculated for the X/MgO/vacuum supercell models. Note that, experimentally, the SOC constant $\alpha_{R}$ can be extracted from angular ‘resolved photoemission (ARPES) Nicolay2001; Reinert2001.
@@ III.3 SOC-induced conductance bottleneck
The orbital symmetries in the electronic structures of V(100) and Fe(100) are mismatched, implying that the heterojunctions based on these materials should exhibit negligible conduction under low-bias conditions. Indeed, early studies on fully epitaxial lateral bcc Fe/V/Fe spin valves have already highlighted the critical role of orbital symmetries in spin-dependent transport across a V/Fe interface Moodera2008.
At the Fermi energy $E_{F}$, (100)V possesses $\Delta_{2}$ orbital symmetry, whereas Fe(100) is dominated by $\Delta_{1}$ symmetry (see Figure 5a). Therefore, MgO acts as an insulating layer that suppresses the $\Delta_{2}$ states and enables TMR in Fe(100)/MgO(100)/Fe(100) structures Parkin2004; Yuasa2004; Butler2001; Mathon2001. Nevertheless, the structural inversion asymmetry in our junctions introduces an interfacial SOC at the V/MgO interface, which works in conjunction with the effective $\Delta_{2}$ filtering barrier of crystalline MgO. As sketched in Figure 5a and discussed in Ref. GonzalezRuano2021-HB-TMR , SOC-driven spin-flip scattering mixes the $\Delta_{2}$ and the $\Delta_{1}$ states, opening a pathway for electron tunneling above $T_{C}$ at low bias, as depicted in Figure 5a.
Another channel contributing (in sequence) to the low bias transport in these epitaxial junctions originates from the so-called “hot spots” in the reciprocal space provided by the $\Delta_{1}$ states of electrons that undergo normal incidence to the barrier, allowing a strong transmission through MgO Butler2001; Mathon2001. The normal-state transport in this system can therefore be approximated by an effective resistance $R_{\text{eq}}=R_{\text{SOC}}+R_{\text{MgO}}$, where the resistance associated with the symmetry-enforced SOC barrier, $R_{\text{SOC}}$, is significantly higher than that the one arising from the conventional MgO barrier, $R_{\text{MgO}}$.
This framework, in which tunneling is governed by symmetry-based spin filtering rather than the barrier height, is supported by our measurements of the low-bias differential conductance $G$ (Figure 5b). Compared with junctions lacking such an interface, $G$ is reduced by roughly two (three) orders of magnitude in the presence of one (two) interfaces. This highlights a unique situation in (superconducting) spintronics in which practically all electrons that are transmitted between V and Fe undergo the effects of SOC.
Figure 5: (a) Transport mechanism in the V/MgO/Fe barrier, sketching the vanadium mixed $\Delta_{2}$ and $\Delta_{1}$ surface states, interfacial spin-orbit coupling and symmetry filtering in the MgO. (b) average conductance for different types of samples of the same lateral size ($20\times 20~\mu\text{m}^{2}$), as a function of the number of V/MgO interfaces for each type.